# Monocular 3D lane detection for Autonomous Driving: Recent Achievements, Challenges, and Outlooks

### v1

@@ IV-A Apollo 3D Lane Synthetic Dataset
ApolloSim [25] - Synthetic (Unity 3D) 1920 × 1080 10.5K 2020 link
OpenLane [12] Waymo [74] Real-world 1920 × 1280 200K 2022 link
ONCE-3DLanes [107] ONCE [61] Real-world 1920 × 1020 211K 2022 link
Synthetic-3D-Lanes [24] - Synthetic (Blender) 360 × 480 306K 2019 -
Real-world 3D-Lanes [24] - Real-world / 85K 2019 -
LLAMAS-3DLD [40] LLAMAS [9] Real-world 1276 × 717 6.6K 2023 -
Pandaset-3DLD [40] Pandaset [104] Real-world 1920 × 1080 6K 2023 -
@@ IV-B OpenLane
OpenLane [12] stands as the first large-scale, real-world 3D lane detection dataset, boasting over 200,000 frames and 880,000 meticulously annotated lanes. Built on the foundation of the influential Waymo Open dataset, OpenLane utilizes the same data format, evaluation pipeline, and a 10Hz sampling rate, courtesy of 64-beam LiDARs, over a 20-second span. This dataset provides exhaustive details for each frame, inclusive of camera intrinsics and extrinsics, along with lane categories, which encompass 14 distinct types such as white dashed lines and road curbsides. Almost 90% of the total lanes are constituted by double yellow solid lanes, and single white solid and dashed lanes. The OpenLane dataset epitomizes a real-world scenario, duly highlighting the long-tail distribution problem. OpenLane includes all lanes in a frame, even accommodating those in the opposite direction, provided there is no dividing curbside. Due to intricate lane topologies like intersections and roundabouts, individual frames can house up to 24 lanes. Approximately 25% of the frames harbor more than six lanes, surpassing the maximum in most current lane datasets. In addition to this, the dataset provides annotations for scene tags such as weather and locations, along with the closest-in-path object (CIPO) - defined as the most pertinent target with respect to the ego vehicle. This ancillary data proves invaluable for subsequent modules in planning and control, and not merely perception. The 3D ground truth of OpenLane is synthesized using LiDAR, thus ensuring high precision and accuracy. The dataset is split between a training set containing 157,000 images and a validation set comprising 39,000 images.
@@ IV-C ONCE-3DLanes
The ONCE-3DLanes dataset [107] manifests as another practical 3D lane detection dataset, meticulously derived from the ONCE autonomous driving repository. This collection encompasses 211,000 visuals captured by a forward-facing camera, complemented by their matching LiDAR point cloud data. Showcasing a wide spectrum of scenes across varied times of day and weather settings, such as sunlit, overcast, and rainy conditions, the dataset casts light on multiple terrains like city centers, residential areas, highways, bridges, and tunnels. This diversity fortifies the dataset as a crucial resource for developing and validating resilient 3D lane detection models under an assortment of real-world scenarios. The dataset is organized into three segments: 3,000 scenes for validation, 8,000 scenes for testing, and the residual 5,000 scenes reserved for training. The training component is supplemented with an extra 200,000 unlabeled scenarios to thoroughly harness the raw data. While the dataset supplies intrinsic camera parameters, it omits extrinsic camera parameters.
@@ IV-D Other Datasets
The paper “3D-LaneNet: End-to-End 3D Multiple Lane Detection” [24] presents two distinct datasets: the Synthetic-3D-Lanes dataset and the 3D-Lanes dataset. The Synthetic-3D-Lanes dataset, created via the open-source graphics engine, Blender, consists of 300K training and 5K testing examples, each featuring a 360×480 pixel image alongside associated ground truth parameters such as 3D lanes, camera height, and pitch. Exhibiting substantial diversity in lane topography, object placement, and scene rendering, this dataset provides a valuable resource for method development and ablation studies. Complementarily, the 3D-Lanes dataset is a real-world ground truth labeled data collection, compiled through the utilization of a multi-sensor setup - a forward-looking camera, a Velodine HDL32 lidar scanner, and a high-precision IMU. This dataset was assembled from six separate drives, each on a different road segment, collectively spanning nearly two hours of driving. With the aid of Lidar and IMU data, aggregated lidar top view images were generated and used alongside a semi-manual annotation tool to establish the ground truth. In total, 85,000 images were annotated, of which 1,000, derived from a separate drive, were designated as the test set, with the remainder serving as the training set. The 3D-Lanes dataset proves instrumental in validating the transferability of the proposed approach to real-world data and enabling qualitative analysis. While the Synthetic-3D-Lanes dataset is made available to the research community, the real-world 3D-Lanes dataset remains proprietary and is not publicly accessible. It’s worth noting that despite its availability, the Synthetic-3D-Lanes dataset has not seen wide adoption for benchmarking in subsequent studies in the field.
@@
The datasets introduced above encompass a variety of publicly available road scenes. Current mainstream research primarily focuses on favorable daytime scenes that are suitable for and conducive to three-dimensional lane detection, with sufficient lighting and favorable weather conditions. Many car companies and original equipment manufacturers in the industry possess a large volume of data; however, they are reluctant to share it publicly due to concerns related to intellectual property, industrial competition, and General Data Protection Regulations (GDPR). As a result, the lack of sufficient annotated data for accurately understanding dynamic weather conditions, such as nighttime, smoggy weather, and edge cases, remains a challenging task in autonomous driving research.
@@ V-A Open Challenges
Although significant research has been conducted by researchers in the field of autonomous driving (AD), and the AD industry is flourishing, there are still open challenges that require researchers’ attention to achieve fully intelligent autonomous driving. These challenges have been individually discussed with support from relevant literature:
@@ V-A1 Coarse-Structured Information
Coarse-Structured Information: Most of the datasets introduced in the literature for 3D lane detection in autonomous driving (AD) are recorded in normal and well-structured infrastructures of advanced cities. The currently developed deep learning models may achieve the best results over structured datasets, but they generalize poorly in many unstructured environments. This aspect of AD demands further attention in terms of data collection, as well as the inclusion of new and effective representation mechanisms in deep learning models.
@@ V-A2 Uncertainty-Aware Decisions
A largely overlooked aspect in lane detection and AD decision-making is the confidence with which models make predictions on input data. However, the confidence of model outputs plays a crucial role in ensuring the safety of autonomous driving. The inherently uncertain nature of the vehicular surroundings seems to have not convinced the community to delve into this matter, as the current methodological trends focus solely on predictive scores. Fortunately, confidence estimation has recently gained attention in the community [5, 63]. Nevertheless, elements from evidential deep learning [71], Bayesian formulations of deep neural networks [38], simpler mechanisms to approximate the output confidence of neural networks (e.g., Monte Carlo dropout [21] or ensembles [44]), and other assorted methods for uncertainty quantification [1] should be progressively incorporated as an additional yet crucial criterion for decision making. When dealing with complex environments, where there is a lack of data that fully represents all possible scenes, the model outputs a significant amount of epistemic uncertainty. Without considering confidence as an additional factor for AD or with current studies focused solely on predictive and/or computational efficiency aspects, there is no guarantee that the emerging 3D lane detection models in the scientific community will be practically useful and transferable to the industry.
@@ V-A3 Weakly supervised Learning Strategies
In the current deep learning-based algorithm models, the majority rely on fully supervised learning strategies, which have high demands for labeled data. This is particularly challenging in the field of 3D lane detection, as general visual sensor data lacks depth information. It is difficult to simply assign 3D information to lanes based on images alone, and alternative sensors such as LiDAR are required to obtain the 3D lane information. This leads to the costly and labor-intensive task of annotating 3D lane data. However, fortunately, both the academic and industrial communities have recognized this issue, and weakly supervised learning strategies have received extensive research and attention in the field of deep learning [reference]. However, in the specific branch of 3D lane detection, there is currently limited research focusing on weakly supervised learning strategies. If we can effectively leverage self-supervised/weakly supervised learning strategies, it would significantly reduce data collection costs and allow for more training data to enhance the performance of lane detection algorithms, thereby further promoting the advancement of the autonomous driving industry.
@@ V-B1 Video-Based 3D Lane Detection for Autonomous Driving
Drawing from the advancements in video-based object detection [102], semantic segmentation [84, 60, 30], and 2D lane detection [34, 112, 111], it’s evident that incorporating video-based techniques significantly improves the precision and dependability of 3D lane detection systems. The core advantage of video-based methods is their capacity to harness temporal data, providing a dynamic view that static images lack. This dynamic perspective is particularly vital in comprehending and predicting complex driving situations in three-dimensional spaces, where the intricacies of lane placement and vehicular interactions intensify. Methods like Recursive Video Lane Detection (RVLD) [34] demonstrate the power of video in capturing ongoing lane shifts and variations over time, a feature immensely beneficial for 3D modeling accuracy. Additionally, the assimilation of video data into these systems allows for an enhanced understanding of spatial dynamics in driving environments, a crucial factor in 3D lane detection where accurate identification of lane positioning and curvature in relation to other road elements is critical. By incorporating sophisticated deep learning techniques used in video-based object detection and semantic segmentation, future iterations of 3D lane detection systems could attain advanced spatial awareness, significantly enhancing the navigation capabilities and safety of autonomous vehicles.
@@ V-B2 Hybrid Approaches and Multimodality
The progression of multi-modal 3D lane detection technologies for autonomous vehicles is greatly accelerated by the incorporation of various sensory inputs, such as cameras, LiDAR, and radar. This integration marks a promising avenue for overcoming the challenges faced by existing camera-reliant systems. This approach, underscored by the successes in multi-modal 3D object detection and semantic segmentation [19], leverages the complementary strengths of each sensor type to enhance detection accuracy and reliability, particularly in challenging environmental and complex driving scenarios [13, 49]. Reflecting on pioneering models such as “Deep Multi-Sensor Lane Detection” [6] and “M^2-3DLaneNet” [54], which have effectively utilized multi-sensor inputs to refine lane boundary estimations and demonstrated robust performance against occlusions and variable lighting conditions, the potential for significant advancements in lane detection technology is evident. The future trajectory in this field should emphasize the exploration of advanced data fusion methods, meticulous sensor calibration, and synchronization techniques, alongside the exploitation of emerging technologies like edge computing for real-time multi-modal data processing.
@@ V-B3 Active Learning and Incremental Learning
Active learning [69] in machine learning refers to the ability of a model to adapt and learn over time and encountering new data during the testing phase and after deployment. In real-world environments, vehicles may encounter randomly occurring unfamiliar scenes and lane topology, which may require AI models to make decisions for further actions such as braking or accelerating to achieve sensible driving operations. Therefore, lane detection techniques should allow for an interactive approach to handle various types of scenes and lane topologies, involving human annotators to label unlabeled data instances and human involvement in the training process [39]. There are different types of active learning techniques, such as membership query synthesis [37], where synthetic data is generated and the parameters of the synthetic data can be adjusted based on the structure of the data [2], which originates from the underlying species of the dataset. On the other hand, the ability of 3D lane detection models to incrementally update their captured knowledge with new data is crucial for their sustainability and continuous improvement. We anticipate that both these capabilities of 3D lane detection models in road understanding will become increasingly important in future research.
@@ V-B4 Adverse Weather Conditions
The progress in developing camera-based 3D lane detection systems for autonomous driving is notably hindered by adverse weather conditions, which significantly impair visibility. Events like heavy rainfall, fog, snow, and dust storms can severely impact the functionality of these systems [113]. The core issue stems from the compromised quality of visual data necessary for the precise detection and segmentation of lane markings, resulting in a decline in reliability and an increase in the likelihood of false negatives or positives. This diminishment in the efficacy of such systems not only elevates safety hazards but also constrains the operational scope of autonomous vehicles. However, recent breakthroughs in object detection and segmentation, as illustrated in research works such as ”ACDC: The Adverse Conditions Dataset with Correspondences for Semantic Driving Scene Understanding” [70] and ”Vehicle Detection and Tracking in Adverse Weather Using a Deep Learning Framework” [27], reveal avenues for enhancing 3D lane detection amidst challenging weather conditions. These studies put forth the advantages of employing deep learning algorithms trained on datasets that include a wide array of adverse weather instances, showcasing the significance of effective data augmentation, domain adaptation tailored to specific conditions, and the employment of semantic segmentation techniques. Employing these approaches, the capabilities of camera-based detection systems can be substantially improved to interpret lane markings accurately and ensure safe navigation even under poor visibility, laying down an optimistic pathway for ongoing research and development in the field of autonomous driving technology.
Figure 9: The test results of the current open-source state-of-the-art LATR [53] algorithm under extreme weather conditions or photography scenarios are depicted in the figure above, with red lines indicating the algorithm’s predicted lane lines and blue lines representing the ground truth lane lines. Figures a and d show scenes with strong exposure during the day and at night, respectively; figures b and e depict rainy scenes during the day and night; and figures c and f show foggy scenes during the day and night. From these test scenarios under harsh conditions, it is evident that the current algorithm performs poorly in detecting lane lines, indicating that there is substantial room for improvement.Additionally, there is a lack of datasets for training and testing under such severe conditions.
@@
The advent of large language models (LLMs) such as ChatGPT [103] has transformed the field of artificial general intelligence (AGI), demonstrating their remarkable zero-shot abilities in tackling a variety of natural language processing (NLP) tasks using customized user prompts or language instructions.
Computer vision encompasses a set of unique challenges and concepts that stand apart from those found in NLP. Vision foundation models typically adhere to a pre-training and subsequent fine-tuning process [92, 14, 73, 93, 18, 80], which, while effective, entails significant additional costs for adaptation to a range of downstream applications. As depicted in Figure 1a, techniques such as multi-task unification [68, 87, 4, 85] aim to endow systems with a broad set of capabilities, yet they often fail to transcend the constraints of predetermined tasks, leaving a noticeable capability deficit in open-ended tasks when juxtaposed with LLMs. The advent of visual prompt tuning [32, 108, 110, 94] provides a novel method for delineating specific vision tasks, like object detection, instance segmentation, and pose estimation, through the use of visual masking. Nonetheless, there is currently no work that combines LLMs with 3D lane line detection. As large language models become more widespread and their capabilities continue to improve, research into LLM-based lane line detection presents an interesting and promising avenue for future exploration.
@@
Balanced Scene 3DLaneNet [24] 53 RTX 2080 89.3 86.4 0.068 0.477 0.015 0.202
Gen-LaneNet [25] 60 RTX 2080 90.1 88.1 0.061 0.496 0.012 0.214
CLGo [52] 75 RTX 2080 94.2 91.9 0.061 0.361 0.029 0.250
PersFormer [12] 21 Tesla V100 - 92.9 0.054 0.356 0.010 0.234
Reconstruct from Top [45] 82 GTX 1080Ti 93.8 91.9 0.049 0.387 0.008 0.213
CurveFormer [7] - - 97.3 95.8 0.078 0.326 0.018 0.219
WS-3D-Lane [3] 26.7 RTX 3080Ti 95.7 93.5 0.027 0.321 0.006 0.215
Anchor3DLane [31] 87.3 RTX 2080Ti 97.2 95.6 0.052 0.306 0.015 0.223
BEV-LaneDet [90] 102 Tesla V100 - 98.7 0.016 0.242 0.020 0.216
LATR [53] 17.75 Tesla A100 97.9 96.8 0.022 0.253 0.007 0.202
D-3DLD [40] 22.6 Tesla A100 95.5 93.3 0.039 0.412 0.011 0.234
Rarely Observed 3DLaneNet [24] 53 RTX 2080 74.6 72.0 0.166 0.855 0.039 0.521
Gen-LaneNet [25] 60 RTX 2080 79.0 78.0 0.139 0.903 0.030 0.539
CLGo [52] 75 RTX 2080 88.3 86.1 0.147 0.735 0.071 0.609
PersFormer [12] 21 Tesla V100 - 87.5 0.107 0.782 0.024 0.602
Reconstruct from Top [45] 82 GTX 1080Ti 85.2 83.7 0.126 0.903 0.023 0.625
CurveFormer [7] - - 97.1 95.6 0.182 0.737 0.039 0.561
WS-3D-Lane [3] 26.7 RTX 3080Ti 88.0 86.0 0.070 0.741 0.015 0.585
Anchor3DLane [31] 87.3 RTX 2080Ti 96.9 94.4 0.094 0.693 0.027 0.579
BEV-LaneDet [90] 102 Tesla V100 - 99.1 0.031 0.594 0.040 0.556
LATR [53] 17.75 Tesla A100 97.3 96.1 0.050 0.600 0.015 0.532
Visual Variations 3D-LaneNet [24] 53 RTX 2080 74.9 72.5 0.115 0.601 0.032 0.230
Gen-LaneNet [25] 60 RTX 2080 87.2 85.3 0.074 0.538 0.015 0.232
CLGo [52] 75 RTX 2080 89.2 87.3 0.084 0.464 0.045 0.312
PersFormer [12] 21 Tesla V100 - 89.6 0.074 0.430 0.015 0.266
Reconstruct from Top [45] 82 GTX 1080Ti 92.1 89.9 0.060 0.446 0.011 0.235
CurveFormer [7] - - 93.0 90.8 0.125 0.410 0.028 0.254
WS-3D-Lane [3] 26.7 RTX 3080Ti 95.8 93.6 0.025 0.338 0.005 0.237
Anchor3DLane [31] 87.3 RTX 2080Ti 93.6 91.4 0.068 0.367 0.020 0.232
BEV-LaneDet [90] 102 Tesla V100 - 96.9 0.027 0.320 0.031 0.256
LATR [53] 17.75 Tesla A100 96.6 95.1 0.045 0.315 0.016 0.228

### v2

@@ IV-A Apollo 3D Lane Synthetic Dataset
ApolloSim [guo2020gen] - Synthetic (Unity 3D) 1920 × 1080 10.5K 2020 link
OpenLane [chen2022persformer] Waymo [Sun_2020_CVPR] Real-world 1920 × 1280 200K 2022 link
ONCE-3DLanes [yan2022once] ONCE [mao2021one] Real-world 1920 × 1020 211K 2022 link
Synthetic-3D-Lanes [garnett20193d] - Synthetic (Blender) 360 × 480 306K 2019 -
Real-world 3D-Lanes [garnett20193d] - Real-world / 85K 2019 -
LLAMAS-3DLD [kim2023d] LLAMAS [9022318] Real-world 1276 × 717 6.6K 2023 -
Pandaset-3DLD [kim2023d] Pandaset [xiao2021pandaset] Real-world 1920 × 1080 6K 2023 -
@@ IV-B OpenLane
OpenLane [chen2022persformer] stands as the first large-scale, real-world 3D lane detection dataset, boasting over 200,000 frames and 880,000 meticulously annotated lanes. Built on the foundation of the influential Waymo Open dataset, OpenLane utilizes the same data format, evaluation pipeline, and a 10Hz sampling rate, courtesy of 64-beam LiDARs, over a 20-second span. This dataset provides exhaustive details for each frame, inclusive of camera intrinsics and extrinsics, along with lane categories, which encompass 14 distinct types such as white dashed lines and road curbsides. Almost 90% of the total lanes are constituted by double yellow solid lanes, and single white solid and dashed lanes. The OpenLane dataset epitomizes a real-world scenario, duly highlighting the long-tail distribution problem. OpenLane includes all lanes in a frame, even accommodating those in the opposite direction, provided there is no dividing curbside. Due to complex lane topologies, such as intersections and roundabouts, a single frame can accommodate up to 24 lanes. Approximately 25% of the frames harbor more than six lanes, surpassing the maximum in most current lane datasets. In addition to this, the dataset provides annotations for scene tags such as weather and locations, along with the closest-in-path object (CIPO) - defined as the most pertinent target with respect to the ego vehicle. This ancillary data proves invaluable for subsequent modules in planning and control, and not merely perception. The 3D ground truth of OpenLane is synthesized using LiDAR, thus ensuring high precision and accuracy. The dataset is split between a training set containing 157,000 images and a validation set comprising 39,000 images.
@@ IV-C ONCE-3DLanes
The ONCE-3DLanes dataset [yan2022once] manifests as another practical 3D lane detection dataset, meticulously derived from the ONCE autonomous driving repository. This collection encompasses 211,000 visuals captured by a forward-facing camera, complemented by their matching LiDAR point cloud data. Showcasing a wide spectrum of scenes across varied times of day and weather settings, such as sunlit, overcast, and rainy conditions, the dataset casts light on multiple terrains like city centers, residential areas, highways, bridges, and tunnels. This diversity fortifies the dataset as a crucial resource for developing and validating resilient 3D lane detection models under an assortment of real-world scenarios. It is organized into three segments: 3,000 scenes for validation, 8,000 scenes for testing, and the residual 5,000 scenes reserved for training. The training component is supplemented with an extra 200,000 unlabeled scenarios to thoroughly harness the raw data. While the dataset supplies intrinsic camera parameters, it omits extrinsic camera parameters.
@@ IV-D Other Datasets
The paper “3D-LaneNet: End-to-End 3D Multiple Lane Detection” [garnett20193d] presents two distinct datasets: the Synthetic-3D-Lanes dataset and the 3D-Lanes dataset. The Synthetic-3D-Lanes dataset, created via the open-source graphics engine, Blender, consists of 300K training and 5K testing examples, each featuring a 360×480 pixel image alongside associated ground truth parameters such as 3D lanes, camera height, and pitch. Exhibiting substantial diversity in lane topography, object placement, and scene rendering, this dataset provides a valuable resource for method development and ablation studies. Complementarily, the 3D-Lanes dataset is a real-world ground truth labeled data collection, compiled through the utilization of a multi-sensor setup - a forward-looking camera, a Velodine HDL32 lidar scanner, and a high-precision IMU. This dataset consists of six independent driving records, each recorded on a different road segment, totaling nearly two hours of driving time. With the aid of Lidar and IMU data, aggregated lidar top view images were generated and used alongside a semi-manual annotation tool to establish the ground truth. In total, 85,000 images were annotated, of which 1,000, derived from a separate drive, were designated as the test set, with the remainder serving as the training set. The 3D-Lanes dataset proves instrumental in validating the transferability of the proposed approach to real-world data and enabling qualitative analysis. While the Synthetic-3D-Lanes dataset is made available to the research community, the real-world 3D-Lanes dataset remains proprietary and is not publicly accessible. It’s worth noting that despite its availability, the Synthetic-3D-Lanes dataset has not seen wide adoption for benchmarking in subsequent studies in the field.
@@
The datasets introduced above encompass a variety of publicly available road scenes. Current mainstream research primarily focuses on favorable daytime scenes that are suitable for and conducive to three-dimensional lane detection, with sufficient lighting and favorable weather conditions. Many car companies and original equipment manufacturers in the industry possess a large volume of data, however, they are reluctant to share it publicly due to concerns related to intellectual property, industrial competition, and General Data Protection Regulations (GDPR). As a result, the lack of sufficient annotated data for accurately understanding dynamic weather conditions, such as nighttime, smoggy weather, and edge cases, remains a challenging task in autonomous driving research.
@@ V-A Open Challenges
Although significant research has been conducted by researchers in the field of AD, and the AD industry is flourishing, there are still open challenges that require researchers’ attention to achieve fully intelligent autonomous driving. These challenges have been individually discussed with support from relevant literature:
@@ V-A1 Coarse-Structured Information
Coarse-Structured Information: Most of the datasets introduced in the literature for 3D lane detection in AD are recorded in normal and well-structured infrastructures of advanced cities. The currently developed deep learning models may achieve the best results over structured datasets, but they generalize poorly in many unstructured environments. This issue of AD demands further attention in terms of data collection, as well as the inclusion of new and effective representation mechanisms in deep learning models.
@@ V-A2 Uncertainty-Aware Decisions
A largely overlooked aspect in lane detection and AD decision-making is the confidence with which models make predictions on input data. However, the confidence of model outputs plays a crucial role in ensuring the safety of autonomous driving. The inherently uncertain nature of the vehicular surroundings seems to have not convinced the community to delve into this matter, as the current methodological trends focus solely on predictive scores. Fortunately, confidence estimation has recently gained attention in the community [arnez2020comparison, michelmore2018evaluating]. Nevertheless, elements from evidential deep learning [sensoy2018evidential], Bayesian formulations of deep neural networks [kendall2017uncertainties], simpler mechanisms to approximate the output confidence of neural networks (e.g., Monte Carlo dropout [gal2016dropout] or ensembles [lakshminarayanan2017simple]), and other assorted methods for uncertainty quantification [abdar2021review] should be progressively incorporated as an additional yet crucial criterion for decision making. When dealing with complex environments, where there is a lack of data that fully represents all possible scenes, the model outputs a significant amount of epistemic uncertainty. Without considering confidence as an additional factor for AD or with current studies focused solely on predictive and/or computational efficiency aspects, there is no guarantee that the emerging 3D lane detection models in the scientific community will be practically useful and transferable to the industry.
@@ V-A3 Weakly supervised Learning Strategies
In current deep learning-based models, the majority rely on fully supervised learning strategies, which have high demands for labeled data. This is particularly challenging in the field of 3D lane detection, as general visual sensor data lacks depth information. It is difficult to simply assign 3D information to lanes based on images alone, and alternative sensors such as LiDAR are required to obtain the 3D lane information. This leads to the costly and labor-intensive task of annotating 3D lane data. Fortunately, both the academic and industrial communities have recognized this issue, and weakly supervised learning strategies have received extensive research and attention in the field of deep learning. However, in the specific branch of 3D lane detection, there is currently limited research focusing on weakly supervised learning strategies. If we can effectively leverage self-supervised/weakly supervised learning strategies, it would significantly reduce data collection costs and allow for more training data to enhance the performance of 3D lane detection algorithms, thereby further advancing the autonomous driving industry.
@@ V-B1 Video-Based 3D Lane Detection for Autonomous Driving
Drawing from the advancements in video-based object detection [wu2019sequence], semantic segmentation [wang2021swiftnet, maninis2018video, hu2018videomatch], and 2D lane detection [Jin2023RecursiveVL, zhang2021vil, zhang2018end], it’s evident that incorporating video-based techniques significantly improves the precision and dependability of 3D lane detection systems. The core advantage of video-based methods is their capacity to harness temporal data, providing a dynamic view that static images lack. This dynamic perspective is particularly vital in comprehending and predicting complex driving situations in three-dimensional spaces, where the intricacies of lane placement and vehicular interactions intensify. Methods like Recursive Video Lane Detection (RVLD) [Jin2023RecursiveVL] demonstrate the power of video in capturing ongoing lane shifts and variations over time, a feature immensely beneficial for 3D modeling accuracy. Additionally, incorporating video data into these systems enhances our understanding of spatial dynamics in driving environments, which is crucial for 3D lane detection. By incorporating sophisticated deep learning techniques used in video-based object detection and semantic segmentation, future iterations of 3D lane detection systems could attain advanced spatial awareness, significantly enhancing the navigation capabilities and safety of autonomous vehicles.
@@ V-B2 Hybrid Approaches and Multimodality
The progression of multi-modal 3D lane detection technologies for autonomous vehicles is greatly accelerated by the incorporation of various sensory inputs, such as cameras, LiDAR, and radar. This integration marks a promising avenue for overcoming the challenges faced by existing camera-reliant systems. This approach, underscored by the successes in multi-modal 3D object detection and semantic segmentation [feng2020deep], leverages the complementary strengths of each sensor type to enhance detection accuracy and reliability, particularly in challenging environmental and complex driving scenarios [chen2019progressive, liang2019multi]. Reflecting on pioneering models such as “Deep Multi-Sensor Lane Detection” [bai2018deep] and “$M^{2}$-3DLaneNet” [luo2022m], which have effectively utilized multi-sensor inputs to refine lane boundary estimations and demonstrated robust performance against occlusions and variable lighting conditions, the potential for significant advancements in lane detection technology is evident. The future trajectory in this field should emphasize the exploration of advanced data fusion methods, meticulous sensor calibration, and synchronization techniques, alongside the exploitation of emerging technologies like edge computing for real-time multi-modal data processing.
@@ V-B3 Active Learning and Incremental Learning
Active learning [ren2021survey] in machine learning refers to the ability of a model to adapt and learn over time and encountering new data during the testing phase and after deployment. In real-world environments, vehicles may encounter randomly occurring unfamiliar scenes and lane topology, which may require AI models to make decisions for further actions such as braking or accelerating to achieve sensible driving operations. Therefore, lane detection techniques should allow for an interactive approach to handle various types of scenes and lane topologies, involving human annotators to label unlabeled data instances and human involvement in the training process [khan2022pmal]. There are different types of active learning techniques, such as membership query synthesis [kelner2020learning], where synthetic data is generated and the parameters of the synthetic data can be adjusted based on the structure of the data [ahmed2022aaqal], which originates from the underlying species of the dataset. On the other hand, the ability of 3D lane detection models to incrementally update their captured knowledge with new data is crucial for their sustainability and continuous improvement. We anticipate that both these capabilities of 3D lane detection models in road understanding will become increasingly important in future research.
@@ V-B4 Adverse Weather Conditions
The progress in developing camera-based 3D lane detection systems for autonomous driving is notably hindered by adverse weather conditions, which significantly impair visibility. Events like heavy rainfall, fog, snow, and dust storms can severely impact the functionality of these systems [zhang2023perception]. The core issue stems from the compromised quality of visual data necessary for the precise detection and segmentation of lane markings, resulting in a decline in reliability and an increase in the likelihood of false negatives or positives. This diminishment in the efficacy of such systems not only elevates safety hazards but also constrains the operational scope of autonomous vehicles. However, recent breakthroughs in object detection and segmentation, as illustrated in research works such as “ACDC: The Adverse Conditions Dataset with Correspondences for Semantic Driving Scene Understanding” [sakaridis2021acdc] and “Vehicle Detection and Tracking in Adverse Weather Using a Deep Learning Framework” [9184996], reveal avenues for enhancing 3D lane detection amidst challenging weather conditions. These studies put forth the advantages of employing deep learning algorithms trained on datasets that include a wide array of adverse weather instances, showcasing the significance of effective data augmentation, domain adaptation tailored to specific conditions, and the employment of semantic segmentation techniques. Employing these approaches, the capabilities of camera-based detection systems can be substantially improved to interpret lane markings accurately and ensure safe navigation even under poor visibility, laying down an optimistic pathway for ongoing research and development in the field of autonomous driving technology.
Figure 9: The test results of the current open-source state-of-the-art LATR [luo2023latr] algorithm under extreme weather conditions or photography scenarios are depicted in the figure above, with red lines indicating the algorithm’s predicted lane lines and blue lines representing the ground truth lane lines. Figures a and d show scenes with strong exposure during the day and at night, respectively; figures b and e depict rainy scenes during the day and night; and figures c and f show foggy scenes during the day and night. From these test scenarios under harsh conditions, it is evident that the current algorithm performs poorly in detecting lane lines, indicating that there is substantial room for improvement.Additionally, there is a lack of datasets for training and testing under such severe conditions.
@@
The advent of large language models (LLMs) such as ChatGPT [wu2023brief] has transformed the field of artificial general intelligence (AGI), demonstrating their remarkable zero-shot abilities in tackling a variety of natural language processing (NLP) tasks using customized user prompts or language instructions.
Computer vision encompasses a set of unique challenges and concepts that stand apart from those found in NLP. Vision foundation models typically adhere to a pre-training and subsequent fine-tuning process [wang2023internimage, chen2022vision, su2023towards, wang2022image, fang2023eva, tao2023siamese], which, while effective, entails significant additional costs for adaptation to a range of downstream applications. Techniques such as multi-task unification [radford2018improving, wang2022ofa, alayrac2022flamingo, wang2022git] aim to endow systems with a broad set of capabilities, yet they often fail to transcend the constraints of predetermined tasks, leaving a noticeable capability deficit in open-ended tasks when juxtaposed with LLMs. The advent of visual prompt tuning [jia2022visual, yao2021cpt, zang2022unified, wang2023images] provides a novel method for delineating specific vision tasks, like object detection, instance segmentation, and pose estimation, through the use of visual masking. Nonetheless, there is currently no work that combines LLMs with 3D lane line detection. As large language models become more widespread and their capabilities continue to improve, research into LLM-based lane line detection presents an interesting and promising avenue for future exploration.
@@
Balanced Scene 3DLaneNet [garnett20193d] 53 RTX 2080 89.3 86.4 0.068 0.477 0.015 0.202
Gen-LaneNet [guo2020gen] 60 RTX 2080 90.1 88.1 0.061 0.496 0.012 0.214
CLGo [liu2022learning] 75 RTX 2080 94.2 91.9 0.061 0.361 0.029 0.250
PersFormer [chen2022persformer] 21 Tesla V100 - 92.9 0.054 0.356 0.010 0.234
Reconstruct from Top [li2022reconstruct] 82 GTX 1080Ti 93.8 91.9 0.049 0.387 0.008 0.213
CurveFormer [bai2022curveformer] - - 97.3 95.8 0.078 0.326 0.018 0.219
WS-3D-Lane [ai2022ws] 26.7 RTX 3080Ti 95.7 93.5 0.027 0.321 0.006 0.215
Anchor3DLane [huang2023anchor3dlane] 87.3 RTX 2080Ti 97.2 95.6 0.052 0.306 0.015 0.223
BEV-LaneDet [wang2023bevlanedet] 102 Tesla V100 - 98.7 0.016 0.242 0.020 0.216
LATR [luo2023latr] 17.75 Tesla A100 97.9 96.8 0.022 0.253 0.007 0.202
D-3DLD [kim2023d] 22.6 Tesla A100 95.5 93.3 0.039 0.412 0.011 0.234
Rarely Observed 3DLaneNet [garnett20193d] 53 RTX 2080 74.6 72.0 0.166 0.855 0.039 0.521
Gen-LaneNet [guo2020gen] 60 RTX 2080 79.0 78.0 0.139 0.903 0.030 0.539
CLGo [liu2022learning] 75 RTX 2080 88.3 86.1 0.147 0.735 0.071 0.609
PersFormer [chen2022persformer] 21 Tesla V100 - 87.5 0.107 0.782 0.024 0.602
Reconstruct from Top [li2022reconstruct] 82 GTX 1080Ti 85.2 83.7 0.126 0.903 0.023 0.625
CurveFormer [bai2022curveformer] - - 97.1 95.6 0.182 0.737 0.039 0.561
WS-3D-Lane [ai2022ws] 26.7 RTX 3080Ti 88.0 86.0 0.070 0.741 0.015 0.585
Anchor3DLane [huang2023anchor3dlane] 87.3 RTX 2080Ti 96.9 94.4 0.094 0.693 0.027 0.579
BEV-LaneDet [wang2023bevlanedet] 102 Tesla V100 - 99.1 0.031 0.594 0.040 0.556
LATR [luo2023latr] 17.75 Tesla A100 97.3 96.1 0.050 0.600 0.015 0.532
Visual Variations 3D-LaneNet [garnett20193d] 53 RTX 2080 74.9 72.5 0.115 0.601 0.032 0.230
Gen-LaneNet [guo2020gen] 60 RTX 2080 87.2 85.3 0.074 0.538 0.015 0.232
CLGo [liu2022learning] 75 RTX 2080 89.2 87.3 0.084 0.464 0.045 0.312
PersFormer [chen2022persformer] 21 Tesla V100 - 89.6 0.074 0.430 0.015 0.266
Reconstruct from Top [li2022reconstruct] 82 GTX 1080Ti 92.1 89.9 0.060 0.446 0.011 0.235
CurveFormer [bai2022curveformer] - - 93.0 90.8 0.125 0.410 0.028 0.254
WS-3D-Lane [ai2022ws] 26.7 RTX 3080Ti 95.8 93.6 0.025 0.338 0.005 0.237
Anchor3DLane [huang2023anchor3dlane] 87.3 RTX 2080Ti 93.6 91.4 0.068 0.367 0.020 0.232
BEV-LaneDet [wang2023bevlanedet] 102 Tesla V100 - 96.9 0.027 0.320 0.031 0.256
LATR [luo2023latr] 17.75 Tesla A100 96.6 95.1 0.045 0.315 0.016 0.228